OpenStudy (anonymous):

Slope 3/4; through (-12,4) y=

OpenStudy (anonymous):

y - 4 = 3(x + 12)/4 y = (3x + 36)/4 + 4 y = (3x + 52)/4

OpenStudy (anonymous):

^ can simplify that to y = 3/4 x + 13